A negative charge of βˆ’0.50 Γ— 10βˆ’6𝐢 exerts a repulsive force of magnitude 0.20 N on an unknown charge 0.20 m away. What is the unknown charge (magnitude and sign)?Β 


1
Expert's answer
2022-03-16T09:43:10-0400
"q2 = - \frac{F * R^2} {k * q1} = - \frac{0.2 * 0.2^2} {9 * 10^9 * 0.5 * 10^{-6}}= - 1.78 * 10^{-6}"